[Hallb-engineering] [Esr-users] [New Logentry] Follow-up Re: Follow-up Re: Hall B Status

Steven Lassiter lassiter at jlab.org
Tue Nov 10 08:09:47 EST 2020

Check with Heidi for replacing CVBI O-rings while still under vacuum.
I believe they have the equipment and procedures for this.
Even if the U-Tube can be vented, it would be helpful to do a trial run under vacuum for experience.

From: Esr-users <esr-users-bounces at jlab.org> on behalf of kashy at jlab.org <kashy at jlab.org>
Sent: Tuesday, November 10, 2020 5:55 AM
To: esr-users at jlab.org <esr-users at jlab.org>; hallb-engineering at jlab.org <hallb-engineering at jlab.org>
Subject: [Esr-users] [New Logentry] Follow-up Re: Follow-up Re: Hall B Status

Follow-up Re: Follow-up Re: Hall B Status<https://logbooks.jlab.org/entry/3857259>

  *   Edit<https://logbooks.jlab.org/entry/3857259/edit?destination=email/send>
  *   Delete<https://logbooks.jlab.org/entry/3857259/delete?destination=email/send>

Lognumber 3857259<https://logbooks.jlab.org/entry/3857259>. Submitted by kashy<https://logbooks.jlab.org/user/kashy> on Tue, 11/10/2020 - 05:51<https://logbooks.jlab.org/entries?start_date=1605001873&end_date=1605009073&book=HBTORUS&book=CLOG>.

Logbooks:       HBTORUS<https://logbooks.jlab.org/book/hbtorus> CLOG<https://logbooks.jlab.org/book/clog>
References:     3857127 - Follow-up Re: Hall B Status<https://logbooks.jlab.org/entry/3857127>

Hall B N2 flow has returned to normal. Continued pumping on the vacuum jacket and leak testing should proceed when the vacuum level is appropriate. This morning I found the heater in the LN2 reservoir in the DBX had a manual position of 30%. This came on because of the reboot of the IOC after it crashed late last week. I set the minimum to 0 and it shut off.

The clean helium flow to the Torus remains at 1.8g/s and is being cooled to 80K and the Torus is now at 80.3K

The Solenoid is at 39.4K and warming at 0.23K/hr. Yesterday Calvin and Morgan leak tested the Solenoid U-tubes. They found the Supply U-tube was leak tight but the Return U-tube had a CVI pump out valve leak through the seal. It was found by pumping on one CVI (no leak found) then switching to the other and the one first pumped on had a leak. The U-tube will be vented today and the "guts" (O-ring, and plug) will be replaced.

The Buffer Dewar is down to 18.3% LL

Next week a team from the Magnet Group is hoping to get in the Hall and Work on the Torus Service Tower. The main task is to change out the ice management heaters on the VCL's
-------------- next part --------------
An HTML attachment was scrubbed...
URL: <https://mailman.jlab.org/pipermail/hallb-engineering/attachments/20201110/20ddbb0c/attachment-0001.html>

More information about the Hallb-engineering mailing list